OpenStudy (anonymous):

A group of bike riders took a 4.0-hour trip. During the first 3.0 hours they traveled a total of 100 miles, but during the last hour they traveled only 20 miles. What was their average speed for the entire trip?

OpenStudy (anonymous):

first speed = 100miles/3h=33(1/3)miles per hour second speed = 20miles/h aver = 1/2(20miles per hour + 33(1/3)miles per hour )

OpenStudy (jamesj):

The average speed is the total distance divided by time, s = d/t. The distance is d = 100 + 20 = 120 miles The time t = 4 hours hence the average speed is s = d/t = 120/4 = 30 miles/hour